OpenStudy (selena2345):

Determine which equation is an identity or whether it has no solution. -8x+14 =-2(4x-7)

OpenStudy (ranga):

Distribute the right hand side and compare it to the left hand side.

OpenStudy (selena2345):

can u show me the work please ? @ranga

OpenStudy (selena2345):

Can u show me the work @ranga

OpenStudy (ranga):

-8x+14 =-2(4x-7) = -2 * 4x + (-2) * (-7) = -8x + 14 The left hand side and right hand side are identical and therefore this equation is an identity. It does not have a UNIQUE solution. But it has infinite number of solution. You can choose any value for x and then solve for y.